OpenStudy (anonymous):

the motion of an object is given parametrically by the equations x=t^2_3t & y=t^3_2t. find all times and locations in which the particle has only vertical and horizantal motion.

OpenStudy (anonymous):

x will have horizontal and the y will have vertical so i believe when x'=0 this will be the horizontal x'=2t-3=0| 2t=3 t=3/2 WHEN t=3/2 X=9/4-9/2=9/4-18/4=-9/2 when y=t^3-2t y'=3t^2-2=0 3t^2=2 t^2=2/3 t=sqrt(2/3) Y(sqrt(2/3))=2/3(sqrt(2/3))-2(sqrt(2/3)) sqrt(2/3)(-4/3)=Y

OpenStudy (anonymous):

X=3/2 Y=-4/2(sqrt(2/3))

OpenStudy (anonymous):

-4/3(sqrt(2/3))=Y